OpenStudy (anonymous):

Find the value of x so that the line passing through (10, 5) and (x, 9) has a slope of -2.

OpenStudy (anonymous):

Use The Law : \[\frac{y_{2}-y_{1}}{x_{2}-x_{1}}=m\] Where m=slope and (10,5)=(x1,y1) ....(x,9)=(x2,y2) Now fill up the Rule,and solve it .. ______ Any questions ?

OpenStudy (anonymous):

hold on ill try solving

OpenStudy (anonymous):

im lost what was the answer

OpenStudy (anonymous):

\[\frac{9-5}{x-10}=\frac{4}{x-10}=-2\]\[-2x+20=4 ,\]\[-2x=-16 , x=8\]

OpenStudy (anonymous):

Any questions ?

OpenStudy (anonymous):

Find the value of y so that so that the line passing through (4, y) and (5, 3) has a slope of 1. on this one what do i do after i get 4/4 ?

OpenStudy (anonymous):

Since the x value increased by one, you need to make the y value increase by 1 in order to have a slope of 1. Therefore, you need to find a number that equals 3 when added to 1, which is 2.

OpenStudy (anonymous):

oh okay thanks